[Brahms-l] Shift report 20060618 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 18 2006 - 00:09:36 EDT
Shifter leader was SB
Shifters were 


<pre>
1. Eight hours of beam in this shift
except for 10-15 minute inturreptions
for polarization measurement

2. Runs 16198 - 16209
   Run 16210 going on

3. No detector problems so far
</pre>
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jun 18 00:33:20 2006

This archive was generated by hypermail 2.1.8 : Sun Jun 18 2006 - 00:33:35 EDT